THE SPECIAL SCHUBERT CALCULUS IS REAL

FRANK SOTTILE

Abstract. We show that the Schubert calculus of enumerative geometry is real, for special Schubert conditions. That is, for any such enumerative problem, there exist real conditions for which all the a priori complex solutions are real.